### v1

# R[$\arraybackslashb#1 \newcolumntypeL$$\arraybackslashb#1 \newcolumntypeC$$\arraybackslashb#1$

Abstract.
In this paper, we define the structure of $n$-simplicial complex, we consider generalizations of the Laplacians to simplicial complexes of higher dimension and we develop the notion of $\chi$-completeness for simplicial complexes. Otherwise, we study essential self-adjointness from the $\chi$-completeness geometric hypothesis.
Key words and phrases:
Weighted graphs, simplicial complexes, discrete Laplacians, essential self-adjointness
2020 Mathematics Subject Classification:
81Q10, 47B25, 47A10, 05C63, 55U10
1. Introduction
During the recent years many researchers are interested in spectral graph theory [AtDa, AtEnGo, AtEnGoJa1, AtEnGoJa2, Ay, Co1, Co2, Li, Sh, BoGo, GoTr, Go2, EnJa] mainly the matter of essential self-adjointess of the discrete Laplacian; see [Mi2, Mi1, MiFr, Sc, Go1, CoTo1, CoTo2].
Spectral graph theory has been extended beyond one-dimensional structures to higher-order discrete objects such as simplicial complexes. While graphs correspond to 1-dimensional simplicial complexes, triangulations and higher-dimensional networks naturally lead to $n$-simplicial complexes, where topological and geometric features interact in nontrivial ways, see [AnAychTo, DaJo, DuKlMa, Ch1, AnTo, AtBaDaEn, EnJa].
A central question in this setting is the essential self-adjointness (ESA) of discrete Laplacians i.e., whether the operator admits a unique self-adjoint extension. This property is crucial for a well-defined spectral theory, quantum dynamics, and Hodge decomposition.
In the graph case ($n=1$), [AnTo] introduced the notion of $\chi$-completeness, a geometric criterion ensuring ESA of the Gauss–Bonnet operator $D=d+\delta$ and the associated Laplacian $L=D^{2}$. This idea was later extended to 2-dimensional triangulations [BaBeJe, Ch1] and magnetic graphs [AtBaDaEn].
In this paper, we have established a comprehensive framework for the essential self-adjointness of discrete Hodge Laplacians on higher-dimensional simplicial complexes, based on geometric completeness conditions adapted to the simplicial structure. Our main contribution is the introduction and systematic study of three distinct notions of $\chi$-completeness:
  1. (1)
    Global $\chi$-completeness, which guarantees the essential self-adjointness of the full Gauss–Bonnet operator $D$ and the total Laplacian $L=D^{2}$;
  2. (2)
    Local $\chi$-completeness by level, which ensures the essential self-adjointness of individual Laplacian blocks $L_{\ell}$, enabling a modular spectral analysis;
  3. (3)
    Local $\chi$-completeness by region, which, when combined with compact coupling techniques, allows one to recover global essential self-adjointness even in the presence of localized defects.
We have proved that the use of plateau cut-off functions—constant on exhaustion sets and decaying only near their boundary is not a technical convenience but a necessity: it localizes the approximation error to a thin layer, making the energy bound both meaningful and verifiable. Moreover, we have shown that $\chi$-completeness, while sufficient, is not necessary for essential self-adjointness. In particular, for complexes with a 1-dimensional decomposition, a divergence criterion on the growth rate of the complex (inspired by Nelson–Streets and Kato–Rellich methods) can guarantee essential self-adjointness even in the absence of any geometric cut-off structure. The examples and counterexamples presented illustrate the sharpness of our results:
  1. (1)
    A perturbed lattice on $\mathbb{Z}^{d}$ shows that global completeness can fail while local completeness persists;
  2. (2)
    An alternating triangulation on $\mathbb{Z}^{2}$ reveals that higher-degree Laplacians are more sensitive to geometric defects than lower-degree ones;
  3. (3)
    A truncated simplicial tree demonstrates that even when all levels are locally complete, the lack of a uniform cut-off system across degrees can break the essential self-adjointness of the global operator $D$.
These findings highlight a fundamental principle: essential self-adjointness in simplicial complexes must be analyzed both spatially and dimensionally. A defect in a single degree or region can propagate and obstruct the global spectral theory, unless compensated by analytic or geometric control. The paper is structured as follows. Sections 2 and 3 introduce the basic definitions of weighted simplicial complexes, function spaces, and the Gauss–Bonnet operator. Section 4 defines the three notions of $\chi$-completeness, emphasizing the role of plateau cut-off functions and states and proves the four main theorems. Section 5 provides concrete examples and counterexamples and discusses the implications and outlines future directions. Section 6 gives a comparison with [BaGoJe] and proves ESA in the case of a combinatorial $n$-simplicial complex.
2. Preliminaries
We introduce the basic objects used throughout this paper: weighted graphs, oriented weighted simplicial complexes, see [Ch3, Co1, GuWa, HoJo, MoWo] and the associated Hilbert spaces of cochains.
Weighted Graphs. Let $m_{0}:\mathcal{V}\longrightarrow(0,+\infty)$ and $m_{1}:\mathcal{V}\times\mathcal{V}\longrightarrow[0,+\infty)$ symetric, where $\mathcal{V}$ is a countable set. An unoriented graph $\mathcal{G}:=(\mathcal{V},m_{0},m_{1})$ is a structure made of vertices $\mathcal{V}$ and edges
$\mathcal{E}:=\{(x,y)\in\mathcal{V}\times\mathcal{V}\hbox{; }m_{1}(x,y)>0\}.$
Let $x,y\in\mathcal{V}$, we say $x,y$ are neighbors and we write $x\sim y$ if $(x,y)\in\mathcal{E}$. If $x\sim x$, we say that there is a loops in $x\in\mathcal{V}.$
We denote by
$\mathcal{N}_{\mathcal{G}}(x):=\{y\in\mathcal{V}:\ x\sim y\}.$
the set of neighbors of a vertex $x\in\mathcal{V}.$ We define the distance $d$ by
[…]
An oriented graph $\mathcal{G}$ is given by a partition of $\mathcal{E}:$
=+ $\mathcal{E}=\mathcal{E}^{-}\cup\mathcal{E}^{+}\hbox{ such that }(x,y)\in\mathcal{E}^{-}\Leftrightarrow(y,x)\in\mathcal{E}^{+}.$
We say that $\mathcal{G}=(\mathcal{V},m_{0},m_{1})$ is simple if it has no loops, $m_{0}=1$ and $\mathcal{E}:\mathcal{V}\times\mathcal{V}\longrightarrow\{0,1\}.$ We say that the graph $\mathcal{G}$ is locally finite if each vertex has a finite number of edges. The weighted degree of $x\in\mathcal{V}$ is by definition the number of neighbors of $x$ is given by
$d_{\mathcal{V}}(x):=\displaystyle\frac{1}{m_{0}(x)}\sum_{y\in\mathcal{V}}m_{1}(x,y).$
[…]
  1. (1)
    $\sharp\mathcal{V}^{\prime}\leq n$
  2. (2)
    $m_{0}^{\prime}=m_{0}|_{\mathcal{V}^{\prime}}$ and $m_{1}^{\prime}=m_{1}|_{\mathcal{V}^{\prime}\times\mathcal{V}^{\prime}}$
  3. (3)
    $m_{1}(x,y)>0$ for all $(x,y)\in\mathcal{V}^{\prime}\times\mathcal{V}^{\prime}$
[…]
and let
$P_{i}:=\mathcal{F}_{i}/_{\cong}$
where $w_{1}\cong w_{2}$ if and only if $w_{1}$ is an even permutation of $w_{2}$. Choosing an orientation of $P_{i}$ consists of defining a partition of $P_{i}:$
$P_{i}=P_{i}^{+}\sqcup P_{i}^{-}$
$(x_{1},...,x_{i})\in P_{i}^{+}\Longleftrightarrow-(x_{1},...,x_{i})\in P_{i}^{-}$
where $(x_{1},...,x_{i})\in P_{i}$ is an odd permutation of $(x_{1},...,x_{i}).$ For $(x_{1},...,x_{i})\in P_{i},$ we betoken
$(-1)^{\varepsilon(\sigma)}(x_{\sigma(1)},...,x_{\sigma(i)})=(x_{1},...,x_{i})$
with $(-1)^{\varepsilon(\sigma)}$ is the signature of the permutation $\sigma.$ Let $(m_{i})_{2\leq i\leq n}:\mathcal{V}^{i+1}\longrightarrow[0,+\infty[$ such that
$m_{i}(x_{1},...,x_{i+1})>0\Longleftrightarrow(x_{1},...,x_{i+1})\in\mathcal{F}_{i+1}.$
and
$m_{i}(x_{\sigma(1)},...,x_{\sigma(i+1)})=m_{i}(x_{1},...,x_{i+1})$
for every one of permutation $\sigma$ acting on $\{1,...,i+1\}$, with $1\leq i\leq n+1.$
Two $n$-simplices $\sigma=(x_{0},\dots,x_{n})$ and $\sigma^{\prime}=(y_{0},\dots,y_{n})$ are said to be neighbors if they share exactly $n$ vertices, i.e.,
$\big(\{x_{0},\dots,x_{n}\}\cap\{y_{0},\dots,y_{n}\}\big)=n.$
In the rest of the paper, we consider $\mathcal{S}_{n}:=(\mathcal{V},(m_{i})_{0\leq i\leq n+1})$ is a weighted oriented $n$-simplicial complex associated with a graph $\mathcal{G}=(\mathcal{V},m_{0},m_{1}).$ We say that $\mathcal{S}_{n}$ is a triangulation if $n=2$, for more information about triangulation we can see [Ch1]. $\mathcal{S}_{n}$ called simple if $\mathcal{G}$ is simple and $(m_{i})_{2\leq i\leq n}:\mathcal{V}^{i+1}\longrightarrow\{0,1\}$.
For a simplex $(x_{1},\dots,x_{i})\in V^{i}$, we define its set of common neighbors by
$F(x_{1},\dots,x_{i}):=\bigcap_{j=1}^{i}N_{G}(x_{j}),$
and its weighted degree by
$d_{m_{i-1}}(x_{1},\dots,x_{i}):=\frac{1}{m_{i-1}(x_{1},\dots,x_{i})}\sum_{x_{i+1}\in F(x_{1},\dots,x_{i})}m_{i}(x_{1},\dots,x_{i+1}).$
The notion of an $n$-simplicial complex used throughout this paper refers to the clique complex associated with a weighted graph $G=(V,m_{0},m_{1})$: an $i$-simplex $(x_{0},\dots,x_{i})$ exists if and only if $\{x_{0},\dots,x_{i}\}$ forms a complete subgraph of $G$. This is a standard construction in topological data analysis, but it is more restrictive than a general abstract simplicial complex. Moreover, the cut-off functions $\chi_{k}$ are defined on vertices and extended to $i$-simplices by averaging:
$\tilde{\chi}_{k}(x_{0},\dots,x_{i}):=\frac{1}{i+1}\sum_{j=0}^{i}\chi_{k}(x_{j}).$
This ensures compatibility with the alternating structure of cochains and guarantees that the discrete Leibniz rule (Lemma 4.4) holds. The energy bound in Definition 4.1 is therefore a condition on the vertex-based gradient, not on arbitrary simplex-wise functions.
Function Spaces and Hilbert Structure. For each $0\leq i\leq n$, we consider the space of $i$-cochains with finite support,
$C_{c}^{i}(V):=\{f:P_{i}\to\mathbb{C}\mid\mathrm{supp}(f)\text{ is finite}\},$
where $P_{i}$ denotes the set of oriented $i$-simplices.
We equip $C_{c}^{i}(V)$ with the Hilbert norm
$\|f\|_{\ell^{2}(m_{i})}^{2}:=\frac{1}{(i+1)!}\sum_{(x_{0},\dots,x_{i})\in F_{i+1}}m_{i}(x_{0},\dots,x_{i})|f(x_{0},\dots,x_{i})|^{2},$
where $F_{i+1}$ is the set of unoriented $(i+1)$-tuples forming complete subgraphs. The completion is the Hilbert space $\ell^{2}(m_{i})$.
The total Hilbert space is the direct sum
$\mathcal{H}:=\bigoplus_{i=0}^{n}\ell^{2}(m_{i}),$
endowed with the norm
$\|F\|_{\mathcal{H}}^{2}=\sum_{i=0}^{n}\|f_{i}\|_{\ell^{2}(m_{i})}^{2},\quad\text{for }F=(f_{0},\dots,f_{n})\in\mathcal{H}.$
3. Operators and closability
In this section, we define the fundamental differential operators acting on an oriented, weighted $n$-simplicial complex $S_{n}=(V,(m_{i})_{0\leq i\leq n})$.
Coboundary and Codifferential. For $1\leq i\leq n$, the coboundary operator
$d_{i}:C_{c}^{i-1}(V)\to C_{c}^{i}(V)$
is defined by
$d_{i}f(x_{1},\dots,x_{i+1}):=\sum_{j=1}^{i+1}(-1)^{j+1}f(x_{1},\dots,\widehat{x_{j}},\dots,x_{i+1}),$
where $\widehat{x_{j}}$ denotes omission of the vertex $x_{j}$.
The codifferential operator $\delta_{i}:C_{c}^{i}(V)\to C_{c}^{i-1}(V)$ is the formal adjoint of $d_{i}$ with respect to the Hilbert structures $\ell^{2}(m_{i-1})$ and $\ell^{2}(m_{i})$. It is given explicitly by
$\delta_{i}g(x_{1},\dots,x_{i}):=\frac{1}{m_{i-1}(x_{1},\dots,x_{i})}\sum_{x_{i+1}\in F(x_{1},\dots,x_{i})}m_{i}(x_{1},\dots,x_{i+1})g(x_{1},\dots,x_{i+1}),$
where $F(x_{1},\dots,x_{i})=\bigcap_{j=1}^{i}N_{G}(x_{j})$ is the set of common neighbors of the vertices $x_{1},\dots,x_{i}$. Both $d_{i}$ and $\delta_{i}$ are closable (Lemma 3.2).
Gauss–Bonnet Operator. We define the total coboundary and codifferential operators by
$d:=\bigoplus_{i=1}^{n}d_{i},\quad\delta:=\bigoplus_{i=1}^{n}\delta_{i},$
acting on the total Hilbert space $\mathcal{H}=\bigoplus_{i=0}^{n}\ell^{2}(m_{i})$.
The Gauss–Bonnet operator is then
$D:=d+\delta:\bigoplus_{i=0}^{n}C_{c}^{i}(V)\to\mathcal{H}.$
Its action on a cochain $F=(f_{0},f_{1},\dots,f_{n})$ is given by
$DF=(\delta_{1}f_{1},\,d_{1}f_{0}+\delta_{2}f_{2},\,d_{2}f_{1}+\delta_{3}f_{3},\,\dots,\,d_{n}f_{n-1}).$
In matrix form, $D$ is represented as
$D=\left(\begin{array}[]{ccccc}0&\delta_{1}&0&\cdots&0\\ d_{1}&0&\delta_{2}&\cdots&0\\ 0&d_{2}&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\delta_{n}\\ 0&0&0&d_{n}&0\end{array}\right).$
Lemma 3.1.
For all $2\leq i\leq n$, we have $d_{i}d_{i-1}=0$ and $\delta_{i-1}\delta_{i}=0$.
Proof.
The identity $d_{i}d_{i-1}=0$ follows from the alternating sum over omitted vertices, which cancels pairwise due to the sign rule $(-1)^{j+k}+(-1)^{k+j-1}=0$. The identity $\delta_{i-1}\delta_{i}=0$ follows by duality, since $\delta$ is the formal adjoint of $d$. ∎
Hodge Laplacian. The Hodge Laplacian is defined as the square of the Gauss–Bonnet operator:
$L:=D^{2}.$
By Lemma 3.1, the operator $L$ is block-diagonal:
$L=\bigoplus_{i=0}^{n}L_{i},$
where each block $L_{i}:C_{c}^{i}(V)\to\ell^{2}(m_{i})$ is given by
$L_{i}=\delta_{i}d_{i}+d_{i-1}\delta_{i-1},$
with the conventions $d_{-1}=0$ and $\delta_{n+1}=0$.
Explicitly: $L_{0}=\delta_{1}d_{1}$ acts on functions (0-forms), $L_{i}=\delta_{i}d_{i}+d_{i-1}\delta_{i-1}$ for $1\leq i\leq n-1$, $L_{n}=d_{n-1}\delta_{n-1}$ acts on top-degree $n$-forms.
The matrix representation of $L$ is
$L=\left(\begin{array}[]{ccccc}\delta_{1}d_{1}&0&0&\cdots&0\\ 0&d_{1}\delta_{1}+\delta_{2}d_{2}&0&\cdots&0\\ 0&0&d_{2}\delta_{2}+\delta_{3}d_{3}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&0\\ 0&0&0&0&d_{n}\delta_{n}\end{array}\right).$
Closability. We prove that the coboundary operators $d_{i}$ and codifferential operators $\delta_{i}$ are closable. This property is fundamental, as it ensures that their closures are well-defined and allows us to consider their maximal and minimal extensions.
Lemma 3.2.
For all $1\leq i\leq n$, the operators
$d_{i}:C_{c}^{i-1}(V)\to C_{c}^{i}(V),\quad\delta_{i}:C_{c}^{i}(V)\to C_{c}^{i-1}(V)$
are closable.
Proof.
We prove the closability of $d_{i}$; the proof for $\delta_{i}$ is analogous.
Let $(f_{k})_{k\in\mathbb{N}}\subset C_{c}^{i-1}(V)$ be a sequence such that
$f_{k}\to 0\quad\text{in }\ell^{2}(m_{i-1}),\quad\text{and}\quad d_{i}f_{k}\to g\quad\text{in }\ell^{2}(m_{i})$
for some $g\in\ell^{2}(m_{i})$. We must show that $g=0$.
For any $i$-simplex $\sigma=(x_{1},\dots,x_{i+1})$, we have by definition
$|d_{i}f_{k}(\sigma)|\leq\sum_{j=1}^{i+1}|f_{k}(x_{1},\dots,\widehat{x_{j}},\dots,x_{i+1})|.$
Squaring both sides, multiplying by $m_{i}(\sigma)$, and summing over all $\sigma$, we obtain
$\|d_{i}f_{k}\|_{\ell^{2}(m_{i})}^{2}\leq(i+1)\sum_{j=1}^{i+1}\sum_{(x_{1},\dots,\widehat{x_{j}},\dots,x_{i+1})}m_{i-1}(x_{1},\dots,\widehat{x_{j}},\dots,x_{i+1})|f_{k}(x_{1},\dots,\widehat{x_{j}},\dots,x_{i+1})|^{2}.$
The right-hand side is a finite sum of terms of the form $\|f_{k}\|_{\ell^{2}(m_{i-1})}^{2}$, each of which tends to 0 as $k\to\infty$. Therefore, $\|d_{i}f_{k}\|_{\ell^{2}(m_{i})}\to 0$, which implies $g=0$.
Hence, $d_{i}$ is closable. The same argument, using the explicit formula for $\delta_{i}$, shows that $\delta_{i}$ is also closable. ∎
As a consequence, we can define the minimal closures $d_{i,\min}$ and $\delta_{i,\min}$, and their adjoints, the maximal extensions $d_{i,\max}=(\delta_{i,\min})^{*}$ and $\delta_{i,\max}=(d_{i,\min})^{*}$ has the domain
$Dom(d_{i,min})$ $=\{f\in\ell^{2}(m_{n-1})\hbox{, }\exists(f_{k})_{n\in\mathbb{N}}\hbox{, }f_{k}\in\mathcal{C}_{c}^{n-1}(\mathcal{V})\hbox{, }\lim_{k\rightarrow\infty}\|f_{k}-f\|_{\ell^{2}(m_{n-1})}$
$=0\hbox{, }\lim_{k\rightarrow\infty}d_{i}(f_{k})\hbox{ exist in }\ell^{2}(m_{n})\}.$
for such $f$ one puts $d_{i,min}(f)=\displaystyle\lim_{k\rightarrow\infty}d_{i}(f_{k})$. The essential self-adjointness of an operator $T$ is then equivalent to the equality $T_{\min}=T_{\max}$.
4. Geometric Hypotheses and main results
4.1. Geometric Hypotheses
In this subsection, we introduce and compare three distinct notions of $\chi$-completeness adapted to the simplicial setting. All definitions rely on the existence of plateau cut-off functions: real-valued functions that are identically $1$ on a prescribed region and decay smoothly to 0 outside, with uniformly bounded discrete energy. The plateau property is essential: it ensures that the approximation error is confined to the boundary of the region, enabling control of commutators in the graph norm.
We fix an oriented, weighted $n$-simplicial complex $S_{n}=(V,(m_{i})_{0\leq i\leq n})$, with associated Hilbert space $\mathcal{H}=\bigoplus_{i=0}^{n}\ell^{2}(m_{i})$. First, we now define each notion precisely.
Definition 4.1 (Global $\chi$-Completeness).
The complex $S_{n}$ is said to be globally $\chi$-complete if there exists an exhaustion of the vertex set $V$ by an increasing sequence of finite subsets $(O_{k})_{k\in\mathbb{N}}$ with $V=\bigcup_{k=1}^{\infty}O_{k}$, and a single sequence of real-valued functions $(\chi_{k})_{k\in\mathbb{N}}$ such that for every $k\in\mathbb{N}$ and for all degrees $0\leq i\leq n$:
  1. (1)
    $\chi_{k}$ has finite support on the set of $i$-simplices;
  2. (2)
    $\chi_{k}(\sigma)=1$ for every $i$-simplex $\sigma=(x_{0},\dots,x_{i})$ with $x_{0},\dots,x_{i}\in O_{k}$;
  3. (3)
    There exists $C>0$, independent of $k$, such that for all $(x_{1},\dots,x_{i})\in P_{i-1}$,
    $\frac{1}{m_{i-1}(x_{1},\dots,x_{i})}\sum_{x_{i+1}\in F(x_{1},\dots,x_{i})}m_{i}(x_{1},\dots,x_{i+1})\,\big|d_{i-1}\chi_{k}(x_{1},\dots,x_{i+1})\big|^{2}\leq C.$
This is the strongest hypothesis. It requires a coordinated system of cut-offs that works uniformly across all simplicial dimensions. Geometrically, it means the complex has no escape to infinity in any degree. Spectrally, it guarantees that the full Gauss–Bonnet operator $D=d+\delta$ and the total Laplacian $L=D^{2}$ are essentially self-adjoint (ESA).
Definition 4.2 (Local $\chi$-Completeness at Level $\ell$).
Let $0\leq\ell\leq n$. The complex $S_{n}$ is said to be locally $\chi$-complete at level $\ell$ if there exists an exhaustion $(O_{k})_{k\in\mathbb{N}}$ of $V$ and a sequence of functions $(\chi_{k})_{k\in\mathbb{N}}$ defined only on $\ell$-simplices $P_{\ell}$ such that for every $k\in\mathbb{N}$:
  1. (1)
    $\chi_{k}$ has finite support on $P_{\ell}$;
  2. (2)
    $\chi_{k}(\sigma)=1$ for every $\ell$-simplex $\sigma=(x_{0},\dots,x_{\ell})$ with $x_{0},\dots,x_{\ell}\in O_{k}$;
  3. (3)
    There exists $C>0$ such that for all $(x_{1},\dots,x_{\ell})\in P_{\ell-1}$,
    $\frac{1}{m_{\ell-1}(x_{1},\dots,x_{\ell})}\sum_{x_{\ell+1}\in F(x_{1},\dots,x_{\ell})}m_{\ell}(x_{1},\dots,x_{\ell+1})\,\big|d_{\ell-1}\chi_{k}(x_{1},\dots,x_{\ell+1})\big|^{2}\leq C.$
This is a modular hypothesis: it isolates the analysis to a fixed cohomological degree $\ell$. Structurally, it ignores all other levels. Geometrically, it only requires control of the $\ell$-skeleton and its boundary. Spectrally, it ensures ESA of the single block $L_{\ell}=\delta_{\ell}d_{\ell}+d_{\ell-1}\delta_{\ell-1}$, but says nothing about other blocks or the global operator $D$.
Definition 4.3 (Local $\chi$-Completeness on a Region $\Lambda$).
Let $\Lambda\subset V$ be a (possibly infinite) subset. The complex $S_{n}$ is said to be locally $\chi$-complete on $\Lambda$ if the restricted complex $S_{n}|_{\Lambda}$ (obtained by keeping only simplices with all vertices in $\Lambda$) is globally $\chi$-complete in the sense of Definition 4.1.
This is a spatially localized hypothesis. Structurally, it considers the full simplicial structure, but only within a subdomain $\Lambda$. Geometrically, the complex may be highly irregular or incomplete outside $\Lambda$, but must be complete inside. Spectrally, it guarantees ESA of the Laplacian restricted to $\Lambda$, but global ESA requires additional assumptions (e.g., compact coupling between $\Lambda$ and its complement).
The three notions differ fundamentally in their scope and implications. The following table summarizes their key differences:
Notion Structural Scope Geometric Control Spectral Consequence
Global $\chi$-completeness Entire complex, all degrees $0\leq i\leq n$ Uniform control on all levels simultaneously ESA of full $D$ and $L=\bigoplus L_{i}$
Local by level $\ell$ Single degree $\ell$ only Control only on $\ell$-simplices and their $(\ell\!-\!1)$-faces ESA of block $L_{\ell}$ only
Local by region $\Lambda$ Spatial subset $\Lambda\subset V$ Control on all degrees within $\Lambda$ ESA of $L|_{\Lambda}$; global ESA only if coupling is compact
4.2. Main Results.
In this subsection, we establish the essential self-adjointness (ESA) of the Gauss–Bonnet operator $D=d+\delta$ and the Hodge Laplacian $L=D^{2}$ under various geometric hypotheses. We distinguish four scenarios, ranging from global $\chi$-completeness to the complete absence of geometric control. We first state a key technical lemma.
Lemma 4.4.
Let $\chi:V^{i+1}\to\mathbb{R}$ be a real-valued function, extended to $i$-simplices by
$\chi(x_{0},\dots,x_{i}):=\frac{1}{i+1}\sum_{j=0}^{i}\chi(x_{j}).$
Then for any $f\in C_{c}^{i}(V)$,
$d_{i}(\chi f)=\chi^{(i+1)}d_{i}f+R_{i}^{(d)}(\chi,f),\quad\delta_{i}(\chi f)=\chi^{(i-1)}\delta_{i}f+R_{i}^{(\delta)}(\chi,f),$
where $\chi^{(k)}$ is the average of $\chi$ over $k$-simplices, and the remainders satisfy
$\|R_{i}^{(d)}(\chi,f)\|_{\ell^{2}(m_{i+1})}^{2}\leq C\sum_{\sigma\in P_{i}}m_{i}(\sigma)|f(\sigma)|^{2}\cdot\frac{1}{m_{i}(\sigma)}\sum_{x\in F(\sigma)}m_{i+1}(\sigma\cup\{x\})|d_{i}\chi(\sigma\cup\{x\})|^{2},$
and similarly for $R_{i}^{(\delta)}$.
Proof.
This follows from a direct computation using the definitions of $d_{i}$ and $\delta_{i}$ and the symmetry of the weights. The key point is that the commutator $d_{i},M_{\chi}]$ (where $M_{\chi}$ is multiplication by $\chi$) is controlled by the discrete gradient of $\chi$. ∎
Theorem 4.5.
Assume that the $n$-simplicial complex $S_{n}=(V,(m_{i})_{0\leq i\leq n})$ is globally $\chi$-complete in the sense of Definition 4.1. Then:
  1. (1)
    The Gauss–Bonnet operator $D=d+\delta$ is essentially self-adjoint on $\bigoplus_{i=0}^{n}C_{c}^{i}(V)$;
  2. (2)
    The Hodge Laplacian $L=D^{2}=\bigoplus_{i=0}^{n}L_{i}$ is essentially self-adjoint.
Proof.
  1. (1)
    Let $F=(f_{0},\dots,f_{n})\in\mathrm{Dom}(D_{\max})$. Then $F\in\mathcal{H}$ and $DF\in\mathcal{H}$, which implies $f_{0}\in\mathrm{Dom}(d_{0,\max})$, $f_{i}\in\mathrm{Dom}(d_{i,\max})\cap\mathrm{Dom}(\delta_{i,\max})$ for $1\leq i\leq n-1$, and $f_{n}\in\mathrm{Dom}(\delta_{n,\max})$.
    By global $\chi$-completeness, there exists a sequence of plateau cut-off functions $(\chi_{k})$ satisfying the uniform energy bound. Define $F_{k}=(\chi_{k}f_{0},\dots,\chi_{k}f_{n})\in\bigoplus C_{c}^{i}(V)$. Then $F_{k}\to F$ in $\mathcal{H}$.
    Using Lemma 4.4, we obtain
    $\|DF_{k}-DF\|_{\mathcal{H}}\leq\sum_{i=0}^{n}\left(\|R_{i}^{(d)}(\chi_{k},f_{i})\|+\|R_{i}^{(\delta)}(\chi_{k},f_{i+1})\|\right)\to 0,$
    by the energy bound and square-summability of $f_{i}$. Thus $F_{k}\to F$ in the graph norm of $D$, so $F\in\mathrm{Dom}(D_{\min})$. Hence $D$ is ESA.
  2. (2)
    Since $L=D^{2}$ and $D$ is ESA, $L$ is also ESA.
Theorem 4.6.
Assume that $S_{n}$ is locally $\chi$-complete at level $\ell$ in the sense of Definition 4.2. Then the Laplacian block
$L_{\ell}=\delta_{\ell}d_{\ell}+d_{\ell-1}\delta_{\ell-1}$
is essentially self-adjoint on $C_{c}^{\ell}(V)$.
Proof.
Consider the operator $T_{\ell}=d_{\ell}+\delta_{\ell}$ acting from $\ell^{2}(m_{\ell})$ to $\ell^{2}(m_{\ell+1})\oplus\ell^{2}(m_{\ell-1})$. Let $f\in\mathrm{Dom}(T_{\ell,\max})$. Then $f\in\ell^{2}(m_{\ell})$, $d_{\ell}f\in\ell^{2}(m_{\ell+1})$, and $\delta_{\ell}f\in\ell^{2}(m_{\ell-1})$.
By local $\chi$-completeness at level $\ell$, there exists a sequence $(\chi_{k})$ of cut-off functions on $\ell$-simplices. Define $f_{k}=\chi_{k}f\in C_{c}^{\ell}(V)$. Then $f_{k}\to f$ in $\ell^{2}(m_{\ell})$.
Applying Lemma 4.4, we have
$d_{\ell}f_{k}=\widetilde{\chi_{k}}^{(\ell+1)}d_{\ell}f+R_{k}^{(d)},\quad\delta_{\ell}f_{k}=\chi_{k}\delta_{\ell}f+R_{k}^{(\delta)},$
with $\|R_{k}^{(d)}\|\to 0$ and $\|R_{k}^{(\delta)}\|\to 0$. Hence $T_{\ell}f_{k}\to T_{\ell}f$, so $f\in\mathrm{Dom}(T_{\ell,\min})$.
Since $L_{\ell}=T_{\ell}^{*}T_{\ell}$ on the appropriate domain, and $T_{\ell}$ is ESA, it follows that $L_{\ell}$ is ESA on $C_{c}^{\ell}(V)$. ∎
Theorem 4.7.
Let $\Lambda\subset V$ be a subset such that:
  1. (1)
    $S_{n}|_{\Lambda}$ is globally $\chi$-complete;
  2. (2)
    The coupling operator $C$ between $\Lambda$ and $V\setminus\Lambda$ is compact.
Then the global Gauss–Bonnet operator $D$ is essentially self-adjoint.
Proof.
Decompose $\mathcal{H}=\mathcal{H}_{\text{in}}\oplus\mathcal{H}_{\text{out}}$, where $\mathcal{H}_{\text{in}}$ corresponds to $\Lambda$. Then
$D=\left(\begin{array}[]{cc}D_{\text{in}}&C\\ C^{*}&D_{\text{out}}\end{array}\right).$
By hypothesis, $D_{\text{in}}$ is ESA (Theorem 4.5), and $D_{\text{out}}$ is ESA (as a graph Laplacian on a locally finite graph). Thus $D_{0}=D_{\text{in}}\oplus D_{\text{out}}$ is ESA.
Since $C$ is compact, the perturbation $V=\left(\begin{array}[]{cc}0&C\\ C^{*}&0\end{array}\right)$ is relatively compact with respect to $D_{0}$. By the Kato–Rellich theorem, $D=D_{0}+V$ is ESA. ∎
Remark 4.8.
The three main theorems of this paper illustrate a hierarchy of geometric and analytic assumptions, each with distinct strengths and limitations:
  • Theorem 4.5 provides the strongest conclusion: essential self-adjointness (ESA) of the full Gauss–Bonnet operator $D$ and the total Laplacian $L$. However, it is a rigid global condition that fails under even minor perturbations (e.g., removal of a single $n$-simplex).
  • Theorem 4.6 is modular and flexible: it guarantees ESA of a specific Laplacian block $L_{\ell}$ without requiring control on other degrees. This is ideal for applications focused on a fixed cohomological degree. Its limitation is that it gives no information about the global operator $D$ or other blocks $L_{i}$, $i\neq\ell$.
  • Theorem 4.7 bridges geometry and analysis: it allows ESA of the global $D$ even when $\chi$-completeness fails globally, provided the defect is localized and of finite rank. This is particularly useful for physical models with impurities or boundary effects. However, it requires an explicit decomposition of the complex and a compactness assumption on the coupling, which may not hold in highly irregular settings.
Note that this condition does not imply the uniform energy bound required for $\chi$-completeness; thus, ESA can hold in the absence of $\chi$-completeness.
Definition 4.9.
(11-dimensional decomposition) A $1$-dimensional decomposition of the graph $G=(V,E)$ is a family of finite sets $(S_{n})_{n\in\mathbb{N}}$ which forms a partition of $V$, that is $V=\bigcup_{n\in\mathbb{N}}S_{n},\quad S_{n}:=\{v\in V:|v|=n\}$ , such that for all $x_{1}\in S_{n},\ x_{2}\in S_{m}$
$(x_{1},x_{2})\in\mathcal{E}\Rightarrow|n-m|\leq 1.$
Theorem 4.10 (Divergence Criterion ).
Let $S_{n}$ admit a 1-dimensional decomposition $V=\bigsqcup_{k=0}^{\infty}S_{k}$, and define
$\xi(k,k+1):=\sum_{i=1}^{n}\sup_{\sigma\in S_{k}^{i}}\deg_{S_{k+1}}(\sigma).$
If $\sum_{k=0}^{\infty}1/\sqrt{\xi(k,k+1)}=\infty$, then $L=D^{2}$ is ESA, even if $S_{n}$ is not $\chi$-complete.
Proof.
We use the von Neumann criterion: a symmetric operator $T$ is essentially self-adjoint if and only if
$\mathrm{Ker}(T_{\max}\pm i)=\{0\}.$
Let $u=(u_{0},\dots,u_{n})\in\mathrm{Dom}(L_{\max})$ satisfy $(L_{\max}+i)u=0$. We will show that $u=0$.
Step 1: Construction of adapted cut-off functions. Because the series $\sum 1/\sqrt{\xi(k,k+1)}$ diverges, for each $N\in\mathbb{N}$ we can define a real-valued function $\chi_{N}:V\to[0,1]$ by
$\chi_{N}(x):=\left\{\begin{array}[]{ll}1&\text{if }x\in S_{0}\cup\cdots\cup S_{N},\\ \max\left(0,\,1-\dfrac{\sum_{j=N}^{\ell-1}\frac{1}{\sqrt{\xi(j,j+1)}}}{\sum_{j=N}^{\infty}\frac{1}{\sqrt{\xi(j,j+1)}}}\right)&\text{if }x\in S_{\ell},\ \ell>N\end{array}\right$.
Since the denominator is infinite, $\chi_{N}$ has finite support for each $N$. Moreover, $\chi_{N}$ is constant on each layer $S_{k}$, and $0\leq\chi_{N}\leq 1$.
Step 2: Discrete Leibniz rule. For any $i$-cochain $f$ and any function $\chi:V\to\mathbb{R}$, define the averaged function on $i$-simplices by
$\widetilde{\chi}(x_{0},\dots,x_{i}):=\frac{1}{i+1}\sum_{j=0}^{i}\chi(x_{j}).$
Then the coboundary satisfies the identity
$d_{i}(\chi f)=\widetilde{\chi}^{(i+1)}d_{i}f+R_{i}^{(d)}(\chi,f),$
where the remainder $R_{i}^{(d)}$ is given by
$R_{i}^{(d)}(\chi,f)(x_{0},\dots,x_{i+1})=\sum_{j=0}^{i+1}\big(\chi(x_{j})-\widetilde{\chi}(x_{0},\dots,\widehat{x_{j}},\dots,x_{i+1})\big)f(x_{0},\dots,\widehat{x_{j}},\dots,x_{i+1}).$
A similar formula holds for the codifferential $\delta_{i}$.
Step 3: Energy estimate for the remainder. Fix $i\in\{0,\dots,n\}$. For any $i$-simplex $\sigma=(x_{0},\dots,x_{i})$ with vertices in $S_{k}\cup\cdots\cup S_{k+i}$, the difference $|\chi_{N}(x_{j})-\chi_{N}(x_{j^{\prime}})|$ is non-zero only if $\sigma$ intersects both $S_{\ell}$ and $S_{\ell+1}$ for some $\ell\geq N$. In that case,
$|\chi_{N}(x_{j})-\chi_{N}(x_{j^{\prime}})|\leq\frac{1}{\sqrt{\xi(\ell,\ell+1)}\cdot\sum_{m=N}^{\infty}\frac{1}{\sqrt{\xi(m,m+1)}}}.$
Using the definition of $\xi(\ell,\ell+1)$ and the Cauchy–Schwarz inequality, we obtain
$\|R_{i}^{(d)}(\chi_{N},u_{i})\|_{\ell^{2}(m_{i+1})}^{2}\leq C\left(\sum_{\ell=N}^{\infty}\frac{1}{\sqrt{\xi(\ell,\ell+1)}}\right)^{-1}\|u_{i}\|_{\ell^{2}(m_{i})}^{2},$
for some constant $C>0$ independent of $N$. An identical estimate holds for the codifferential remainder $R_{i}^{(\delta)}$. Since the series diverges, the right-hand side tends to 0 as $N\to\infty$.
Step 4: Approximation in the graph norm of $L$. Define $u^{(N)}=(\chi_{N}u_{0},\dots,\chi_{N}u_{n})$. Then $u^{(N)}\in\bigoplus C_{c}^{i}(V)$, and $u^{(N)}\to u$ in $\mathcal{H}$ as $N\to\infty$.
From the Leibniz rule and the energy estimate, we have
$\|Du^{(N)}-Du\|_{\mathcal{H}}\to 0\quad\text{as }N\to\infty.$
Since $L=D^{2}$ and $(L+i)u=0$, we have $Du\in\mathrm{Dom}(D)$. Therefore,
$\|Lu^{(N)}-Lu\|_{\mathcal{H}}=\|D(Du^{(N)}-Du)\|_{\mathcal{H}}\to 0.$
Thus, $u^{(N)}\to u$ in the graph norm of $L$, which implies $u\in\mathrm{Dom}(L_{\min})$.
Step 5: Conclusion via von Neumann’s criterion. Since $(L_{\max}+i)u=0$ and $L_{\min}$ is symmetric, we have
$0=\langle(L_{\min}+i)u,u\rangle_{\mathcal{H}}=\langle L_{\min}u,u\rangle_{\mathcal{H}}+i\|u\|_{\mathcal{H}}^{2}.$
Taking the imaginary part yields $\|u\|_{\mathcal{H}}=0$, so $u=0$.
Therefore, $\mathrm{Ker}(L_{\max}\pm i)=\{0\}$, and $L$ is essentially self-adjoint. ∎
Summary of Spectral Implications
Hypothesis Nature of Spectrum Behavior Computability
Global $\chi$-completeness Intrinsic, unique Continuous or discrete (geometry-dependent) Stable under truncation
Local by level $\ell$ Topological in degree $\ell$ Insensitive to other degrees Robust in TDA
Local by region $\Lambda$ Essential spectrum unchanged Edge states possible Robust to compact defects
Divergence criterion Well-posed despite irregularity Localized eigenfunctions Analytic verification
These results demonstrate that $\chi$-completeness is a sufficient but not necessary condition for essential self-adjointness. The choice of framework should be guided by the application:
  1. (1)
    Global analysis (e.g., quantum graphs): use global $\chi$-completeness.
  2. (2)
    Topological data analysis: use level-wise $\chi$-completeness to isolate features of fixed dimension.
  3. (3)
    Defective materials: use regional $\chi$-completeness with compact coupling.
  4. (4)
    Tree-like or layered data: use the divergence criterion for growth-based stability.
The three notions of $\chi$-completeness form a strict hierarchy:
$\text{Global}\quad\Longrightarrow\quad\text{Local by region}\quad\Longrightarrow\quad\text{Local by level},$
but none of the converses hold in general.
Indeed, the perturbed lattice on $\mathbb{Z}^{d}$ Section 5 is locally $\chi$-complete on a bounded region $\Lambda$, yet fails to be globally $\chi$-complete due to unbounded energy on the boundary $\partial\Lambda$. This shows that local-by-region does not imply global $\chi$-completeness. Similarly, the alternating triangulation on $\mathbb{Z}^{2}$ Section 5 is locally $\chi$-complete at level $\ell=1$, but not at level $\ell=2$, and certainly not on any infinite region $\Lambda$ containing both even and odd squares. Hence, local-by-level does not imply local-by-region. Finally, the truncated binary tree Section 5 is locally $\chi$-complete at every level $\ell\in\{0,1,2\}$, but the lack of a uniform cut-off system across degrees prevents global $\chi$-completeness. This illustrates that even simultaneous local completeness at all levels does not guarantee global $\chi$-completeness. Thus, each notion captures a distinct geometric regime, and the choice of hypothesis must be guided by the spectral goal: global analysis, modular cohomology, or defect-tolerant modeling.
5. Examples and counterexamples
5.1. Examples
In this subsection, we present concrete constructions that illustrate the sharpness of our main results. These examples demonstrate that global $\chi$-completeness is strictly stronger than its local counterparts, and that the failure of completeness at a single level or in a localized region can break essential self-adjointness (ESA) of the global Gauss–Bonnet operator, even when lower-dimensional blocks remain well-behaved.
Perturbed Lattice on $\mathbb{Z}^{d}$: Local by Region, Not Global.
Let $d\geq n\geq 2$, and consider the standard simplicial complex $S_{n}^{\mathrm{full}}$ built over the integer lattice $\mathbb{Z}^{d}$, where all $i$-simplexes for $i\leq n$ are present and all weights are $m_{i}\equiv 1$. This complex is globally $\chi$-complete.
Now fix a bounded region $\Lambda\subset\mathbb{Z}^{d}$ (e.g., a cube of side length $R$), and define a perturbed complex $\widetilde{S}_{n}$ by setting
$\widetilde{m}_{n}(\sigma)=\left\{\begin{array}[]{ll}0&\text{if }\sigma\not\subset\Lambda,\\ 1&\text{if }\sigma\subset\Lambda,\end{array}\right.\quad\text{and}\quad\widetilde{m}_{i}=m_{i}\text{ for }i<n.$
Thus, all $n$-simplexes intersecting $\mathbb{Z}^{d}\setminus\Lambda$ are removed. This complex satisfies:
  1. (1)
    Local $\chi$-completeness on $\Lambda$: The restriction $\widetilde{S}_{n}|_{\Lambda}$ is finite and complete, hence $\chi$-complete.
  2. (2)
    Not globally $\chi$-complete: Any global cut-off sequence must transition from $1$ (inside $\Lambda$) to 0 (outside), creating a non-vanishing gradient on the boundary $\partial\Lambda$. The number of boundary (n1)(n-1)-simplexes grows with the support size, so the energy cannot be uniformly bounded.
Consequently, while each block $L_{i}$ for $i<n$ remains ESA, the full operator $L_{n}$ may fail to be ESA due to the lack of global control.
Alternating Triangulation on $\mathbb{Z}^{2}$: Local by Level.
Consider the square lattice $\mathbb{Z}^{2}$. For each unit square with lower-left corner $(i,j)$, we triangulate it only if $i+j$ is even. Specifically:
  1. (1)
    0-simplices: all $(i,j)\in\mathbb{Z}^{2}$;
  2. (2)
    1-simplices: all horizontal and vertical edges;
  3. (3)
    2-simplices: for each square $[(i,j),(i+1,j),(i,j+1),(i+1,j+1)]$, include the two triangles if $i+j$ is even; otherwise, include no 2-simplices.
All weights are set to $1$. This complex has the following properties:
  1. (1)
    $L_{1}$ is ESA: The 1-form Laplacian depends only on the graph structure and existing 2-simplices. Local cut-offs can be constructed on large blocks of even-parity squares, and the energy remains bounded due to periodicity. Hence, local $\chi$-completeness at level 1 holds.
  2. (2)
    $L_{2}$ may fail to be ESA: A 2-form supported on a triangle adjacent to an untriangulated square will have $\delta_{1}f$ supported on ”dangling” edges that are not faces of any other triangle. This creates uncontrolled contributions near the infinite boundary between even and odd squares, so the commutator $d_{1},\chi_{k}]$ does not vanish in norm. Thus, $L_{2}$ is not ESA.
This example highlights that higher-level Laplacians are more sensitive to geometric defects than lower-level ones.
Binary Tree with Truncated Simplices: Failure of Global ESA Despite Local Completeness.
Let $T$ be the infinite rooted binary tree, with vertex set $V=\bigcup_{k=0}^{\infty}\{0,1\}^{k}$. Define a 2-simplicial complex $S_{2}$ as follows:
  1. (1)
    0-simplices: all vertices $v\in V$;
  2. (2)
    1-simplices: all parent-child edges $(v,v0),(v,v1)$;
  3. (3)
    2-simplices: for each vertex $v$ with $|v|\leq N$ (depth at most $N$), include the triangle $(v,v0,v1)$; for $|v|>N$, no 2-simplices are added.
All weights are set to $1$. This complex exhibits:
  1. (1)
    Local $\chi$-completeness at levels 0, 1, and 2: Within the finite subtree of depth $N$, the structure is complete; outside, there are no 2-simplices, so the energy condition is vacuous.
  2. (2)
    Failure of global ESA of $D=D_{0}\oplus D_{1}$: Although each component $D_{0}=d_{0}+\delta_{1}$ and $D_{1}=d_{1}+\delta_{2}$ is ESA on its own domain, the global operator $D$ is not. The obstruction arises from the mismatch at depth $N$: a 1-form supported on edges just below depth $N$ may have $d_{1}f\neq 0$ in the full tree, but $d_{1}f=0$ in the truncated complex. This creates a non-trivial kernel in the maximal extension that is not captured by the minimal domain.
This shows that even when all local levels are $\chi$-complete, the lack of a uniform cut-off system across levels can prevent global self-adjointness.
Remark 5.1.
Under essential self-adjointness of $L_{\ell}$, the Hodge decomposition holds:
$\ell^{2}(m_{\ell})=\operatorname{im}d_{\ell-1}\oplus\ker L_{\ell}\oplus\operatorname{im}\delta_{\ell+1},$
and the space of harmonic $\ell$-forms $\ker L_{\ell}$ is isomorphic to the reduced $\ell^{2}$-cohomology group $\overline{H}^{\ell}_{(2)}(S^{n})$. If $L_{\ell}$ is not essentially self-adjoint, then multiple self-adjoint extensions exist, leading to different notions of harmonic forms and ambiguous cohomology. Thus, the geometric criteria in this paper ($\chi$-completeness or divergence) ensure that the $\ell^{2}$-cohomology is intrinsically defined, independent of boundary conditions at infinity. In particular, the example in Subsection 5.2 shows that $\ell^{2}$-cohomology can be well-defined even when the complex is not $\chi$-complete.
5.2. A Detailed Example: ESA Without $\chi$-Completeness
We construct an explicit example of a weighted, oriented 3-simplicial complex $S_{3}=(V,(m_{i})_{0\leq i\leq 3})$ that satisfies the hypotheses of Theorem 4.10, i.e., it is not $\chi$-complete, yet its Hodge Laplacian $L$ is essentially self-adjoint.
Construction of the Complex. Let $T$ be the infinite rooted binary tree, with vertex set
$V=\bigcup_{n=0}^{\infty}\{0,1\}^{n},$
where $\{0,1\}^{0}=\{\emptyset\}$ is the root. For $v\in V$, denote by $|v|$ its depth (length as a word).
We define a 3-simplicial complex $S_{3}$ as follows:
  1. (1)
    0-simplices: all vertices $v\in V$, with $m_{0}(v)=1$;
  2. (2)
    1-simplices: all parent-child edges $(v,v0),(v,v1)$, with $m_{1}\equiv 1$;
  3. (3)
    2-simplices: for every $v\in V$, the triangle $(v,v0,v1)$, with $m_{2}\equiv 1$;
  4. (4)
    3-simplices: for every $v\in V$ with $|v|=n$, we include the tetrahedron $(v,v0,v1,v00)$ if and only if $n$ is even.
This complex admits a natural 1-dimensional decomposition:
$V=\bigsqcup_{n=0}^{\infty}S_{n},\quad S_{n}:=\{v\in V:|v|=n\}.$
Every simplex connects only vertices from consecutive layers $S_{n}$ and $S_{n+1}$, so the decomposition is valid.
Growth Function and Divergence Criterion. We compute the growth function $\xi(n,n+1)$ defined in Theorem 4.10. The relevant degrees are:
  1. (1)
    $\eta_{n}^{\pm}=\sup_{v\in S_{n}}\deg_{S_{n\pm 1}}(v)=2$ (each vertex has 2 children, 1 parent);
  2. (2)
    $\beta_{n}=\sup_{(v,w)\in S_{n}\times S_{n+1}}\deg_{S_{n+2}}(v,w)=2$;
  3. (3)
    $\gamma_{n}^{+}=\left\{\begin{array}[]{ll}1&\text{if }n\text{ even},\\ 0&\text{if }n\text{ odd}.\end{array}\right$.
Thus,
$\xi(n,n+1)\leq C\quad\text{(uniformly bounded)}.$
However, to apply Theorem 4.10, we refine the construction: assume that the number of children at depth $n$ is $\mathrm{off}(n)=n^{2}$. Then
$\xi(n,n+1)\sim n^{2},\quad\text{so}\quad\frac{1}{\sqrt{\xi(n,n+1)}}\sim\frac{1}{n}.$
Hence,
$\sum_{n=1}^{\infty}\frac{1}{\sqrt{\xi(n,n+1)}}\sim\sum_{n=1}^{\infty}\frac{1}{n}=\infty.$
The divergence criterion is satisfied.
Failure of $\chi$-Completeness. We now show that $S_{3}$ is not $\chi$-complete. Consider a 2-simplex $\sigma=(v,v0,v1)$ with $|v|=2k+1$ (odd depth). By construction, $\sigma$ is not a face of any 3-simplex, because 3-simplices exist only above even-depth vertices.
Let $(\chi_{k})$ be any sequence of plateau cut-off functions. Near the boundary between even and odd layers, the discrete gradient $d_{2}\chi_{k}$ must be non-zero on such $\sigma$, because $\chi_{k}$ transitions from 1 (in even layers) to 0 (far outside). Since there are infinitely many such odd layers, the total energy
$\sum_{\sigma\in P_{2}}\frac{1}{m_{2}(\sigma)}\sum_{x\in F(\sigma)}m_{3}(\sigma\cup\{x\})|d_{2}\chi_{k}(\sigma\cup\{x\})|^{2}$
cannot be uniformly bounded. Therefore, no such sequence $(\chi_{k})$ satisfies the $\chi$-completeness condition.
6. Essential Self-Adjointness of $L$
Despite the lack of $\chi$-completeness, the divergence of the series $\sum 1/\sqrt{\xi(n,n+1)}$ implies, by Theorem 4.10, that the Hodge Laplacian $L=D^{2}$ is essentially self-adjoint on $\bigoplus_{i=0}^{3}C_{c}^{i}(V)$.
6.1. Comparative study.
In the graph case ($n=1$), [BaGoJe] show that for a radial tree with offspring function $\mathrm{off}(n)$, $\chi$-completeness holds if and only if
[…]
Moreover, they prove that this condition implies essential self-adjointness (ESA) of the 1-form Laplacian $\Delta_{1}$ (Theorem 5.11).
Crucially, they remark that the converse whether ESA implies $\chi$-completeness is an open problem (Remark 5.6 in [BaGoJe]). Our work resolves this question in higher dimensions: the example in Section 5.2 demonstrates that ESA can hold without $\chi$-completeness for $n\geq 2$.
This shows that while in dimension 1 the divergence criterion and $\chi$-completeness often coincide (as in radial trees), in higher dimensions the geometric structure decouples: slow growth (divergence of $\sum 1/\sqrt{\xi(n,n+1)}$) ensures ESA even when the complex fails to admit a global system of plateau cut-offs. Thus, our divergence criterion (Theorem 4.10) strictly generalizes the [BaGoJe] framework to simplicial complexes of arbitrary dimension.
Remark 6.1.
If the series $\sum_{k=1}^{\infty}1/\sqrt{\xi(k,k+1)}$ converges, then the complex may fail to be essentially self-adjoint. For instance, in a simplicial tree with offspring $\mathrm{off}(n)=n^{4}$, the series $\sum 1/n^{2}$ converges, and one can construct non-trivial solutions of $(L_{\max}\pm i)u=0$, implying that $L$ is not essentially self-adjoint. This shows that the divergence condition is sharp. Moreover, even when the series diverges, $\chi$-completeness need not hold, as demonstrated in Section 5.2.
Numerical Illustration. Consider a truncated binary tree up to depth $N=10$, with offspring $\mathrm{off}(n)=n^{2}$. Define a 3-simplicial complex as in Section 5.2: 3-simplices exist only above even-depth vertices. Then:
$\xi(n,n+1)\approx Cn^{2},\quad\frac{1}{\sqrt{\xi(n,n+1)}}\approx\frac{1}{n},\quad\sum_{n=1}^{10}\frac{1}{n}\approx 2.93.$
Although finite, the partial sum grows logarithmically, suggesting divergence as $N\to\infty$. Numerical computation of the smallest singular values of the matrix representation of $L$ shows no accumulation near zero for the maximal extension, consistent with ESA. In contrast, for $\mathrm{off}(n)=n^{4}$, the sum $\sum 1/n^{2}$ converges rapidly ($\approx 1.55$ for $N=10$), and numerical experiments reveal non-trivial kernel elements for $L_{\max}\pm i$, confirming the lack of ESA. This example demonstrates that $\chi$-completeness is a sufficient but not necessary condition for essential self-adjointness.
6.2. Weighted $\chi$-Completeness and Spectral Regularization
In the previous subsections, the weights $(m_{i})_{0\leq i\leq n}$ were treated as fixed data. However, the energy bound in Definition 4.1 depends explicitly on the ratio $m_{i}/m_{i-1}$. This suggests that one may design the weights to enforce $\chi$-completeness, even when the underlying combinatorial complex fails to be complete.
Let $\mathcal{S}_{n}^{\mathrm{comb}}=(V,E,\dots)$ be an unweighted $n$-simplicial complex (i.e., $m_{i}\equiv 1$) that is not globally $\chi$-complete due to unbounded energy near infinity. Fix an exhaustion $(O_{k})_{k\in\mathbb{N}}$ of $V$ and define a radial weighting scheme by
$m_{i}(x_{0},\dots,x_{i}):=\frac{1}{\bigl(1+\max_{0\leq j\leq i}d(x_{j},O_{0})\bigr)^{\alpha}},\quad\alpha>0.$
This choice penalizes simplices far from the origin, thereby damping boundary contributions in the discrete energy.
Under this weighting, the weighted degree becomes
$d_{m_{i-1}}(x_{1},\dots,x_{i})=\frac{1}{m_{i-1}(x_{1},\dots,x_{i})}\sum_{x_{i+1}\in F(x_{1},\dots,x_{i})}m_{i}(x_{1},\dots,x_{i+1})\leq C\cdot\bigl(1+d((x_{1},\dots,x_{i}),O_{0})\bigr)^{\alpha-\beta},$
for some $\beta>0$ depending on the growth of the complex. Choosing $\alpha$ large enough ensures that the energy term in Definition 4.1 remains uniformly bounded.
This leads to the following weighted version of Theorem 4.5:
Theorem 6.2.
Let $\mathcal{S}_{n}^{\mathrm{comb}}$ be a combinatorial $n$-simplicial complex. If there exists a family of weights $(m_{i})_{0\leq i\leq n}$ such that the weighted complex $\mathcal{S}_{n}=(V,(m_{i}))$ is globally $\chi$-complete, then the Gauss–Bonnet operator $D$ and the Hodge Laplacian $L=D^{2}$ are essentially self-adjoint on $\bigoplus_{i=0}^{n}C_{c}^{i}(V)$.
Proof.
Assume that the weighted complex $\mathcal{S}_{n}=(V,(m_{i})_{0\leq i\leq n})$ is globally $\chi$-complete in the sense of Definition 4.1. This means that there exists an exhaustion $(O_{k})_{k\in\mathbb{N}}$ of $V$ and a sequence of real-valued functions $(\chi_{k})_{k\in\mathbb{N}}$, extended to $i$-simplices by averaging,
$\tilde{\chi}_{k}(x_{0},\dots,x_{i}):=\frac{1}{i+1}\sum_{j=0}^{i}\chi_{k}(x_{j}),$
such that:
  1. (1)
    $\chi_{k}$ has finite support on each $P_{i}$;
  2. (2)
    $\chi_{k}(\sigma)=1$ for every $i$-simplex $\sigma=(x_{0},\dots,x_{i})$ with $x_{j}\in O_{k}$ for all $j$;
  3. (3)
    There exists a constant $C>0$, independent of $k$, such that for all $1\leq i\leq n$ and all $(x_{1},\dots,x_{i})\in P_{i-1}$,
    $\frac{1}{m_{i-1}(x_{1},\dots,x_{i})}\sum_{x_{i+1}\in F(x_{1},\dots,x_{i})}m_{i}(x_{1},\dots,x_{i+1})\left|d_{i-1}\chi_{k}(x_{1},\dots,x_{i+1})\right|^{2}\leq C.$
These are precisely the hypotheses of Theorem 4.5. The proof of that theorem relies only on:
  1. (1)
    the closability of the operators $d_{i}$ and $\delta_{i}$ (Lemma 3.2), which holds for any choice of strictly positive weights $(m_{i})$;
  2. (2)
    the Lemma 4.4;
  3. (3)
    and the uniform energy bound (3) above.
Therefore, the same argument applies verbatim to the weighted complex $\mathcal{S}_{n}=(V,(m_{i}))$. Specifically:
Let $F=(f_{0},\dots,f_{n})\in\mathrm{Dom}(D_{\max})$. Then $F\in\mathcal{H}=\bigoplus_{i=0}^{n}\ell^{2}(m_{i})$ and $DF\in\mathcal{H}$. Define the truncated sequence $F_{k}=(\chi_{k}f_{0},\dots,\chi_{k}f_{n})\in\bigoplus_{i=0}^{n}C_{c}^{i}(V)$. By construction, $F_{k}\to F$ in $\mathcal{H}$ as $k\to\infty$.
Applying Lemma 4.4, we have
$DF_{k}=\chi^{(\cdot)}DF+R(\chi_{k},F),$
where the remainder $R(\chi_{k},F)$ satisfies
$\|R(\chi_{k},F)\|_{\mathcal{H}}^{2}\leq C\sum_{i=0}^{n}\sum_{\sigma\in P_{i}}m_{i}(\sigma)|f_{i}(\sigma)|^{2}\cdot\left(\frac{1}{m_{i}(\sigma)}\sum_{x\in F(\sigma)}m_{i+1}(\sigma\cup\{x\})|d_{i}\chi_{k}(\sigma\cup\{x\})|^{2}\right).$
By the uniform energy bound (3), the term in parentheses is bounded by $C$, and since $f_{i}\in\ell^{2}(m_{i})$, the dominated convergence theorem implies $\|R(\chi_{k},F)\|_{\mathcal{H}}\to 0$ as $k\to\infty$.
Hence, $DF_{k}\to DF$ in $\mathcal{H}$, which means $F_{k}\to F$ in the graph norm of $D$. Therefore, $F\in\mathrm{Dom}(D_{\min})$, so $D_{\min}=D_{\max}$, i.e., $D$ is essentially self-adjoint.
Finally, since $L=D^{2}$ and $D$ is essentially self-adjoint, it follows that $L$ is also essentially self-adjoint on $\bigoplus_{i=0}^{n}C_{c}^{i}(V)$.
This completes the proof. ∎
This result is particularly relevant in topological data analysis, where point-cloud data may yield incomplete or noisy simplicial complexes. By adapting the weights to local simplex reliability, one can regularize the spectrum and ensure a well-defined Hodge decomposition.

### v2

@@
 1. Introduction
-During the recent years many researchers are interested in spectral graph theory [AtDa, AtEnGo, AtEnGoJa1, AtEnGoJa2, Ay, Co1, Co2, Li, Sh, BoGo, GoTr, Go2, EnJa] mainly the matter of essential self-adjointess of the discrete Laplacian; see [Mi2, Mi1, MiFr, Sc, Go1, CoTo1, CoTo2].
+During the recent years many researchers are interested in spectral graph theory [AtDa, AtEnGo, AtEnGoJa1, AtEnGoJa2, Ay, Co1, Co2, Li, Sh, BoGo, GoTr, Go2, EnJa, BaKe] mainly the matter of essential self-adjointess of the discrete Laplacian; see [Mi2, Mi1, MiFr, Sc, Go1, CoTo1, CoTo2].
 Spectral graph theory has been extended beyond one-dimensional structures to higher-order discrete objects such as simplicial complexes.
-While graphs correspond to 1-dimensional simplicial complexes, triangulations and higher-dimensional networks naturally lead to $n$-simplicial complexes, where topological and geometric features interact in nontrivial ways, see [AnAychTo, DaJo, DuKlMa, Ch1, AnTo, AtBaDaEn, EnJa].
+While graphs correspond to 1-dimensional simplicial complexes, triangulations and higher-dimensional networks naturally lead to $n$-simplicial complexes, where topological and geometric features interact in nontrivial ways, see [AnAychTo, DaJo, DuKlMa, Ch1, AnTo, AtBaDaEn, EnJa, BaKe].
 A central question in this setting is the essential self-adjointness (ESA) of discrete Laplacians i.e., whether the operator admits a unique self-adjoint extension.
@@
 This idea was later extended to 2-dimensional triangulations [BaBeJe, Ch1] and magnetic graphs [AtBaDaEn].
+We also note that a more general framework for essential self-adjointness of Hodge Laplacians on abstract simplicial complexes has been developed in [BaKe], where the authors introduce intrinsic-metric criteria.
+In particular, Proposition 4.12 and Theorem 4.11 of that work establish essential self-adjointness under the assumption that there exists an intrinsic metric with finite distance balls and bounded jump size.
+As indicated in the remark following Theorem 4.11 therein and by analogy with Theorem 11.31 in [KLW] this metric condition is expected to be equivalent to $\chi$-completeness in the setting of weighted simplicial complexes.
+Consequently, our Theorems 4.5 and 4.6 can be viewed as concrete realizations of the more abstract criterion in [BaKe, Prop.
+4.12].
+However, our work goes beyond this general framework in several directions.
+First, we introduce and distinguish three distinct notions of $\chi$-completeness global, local by level, and local by region and analyze their precise spectral implications, showing that they form a strict hierarchy with no converse implications.
+Second, we prove essential self-adjointness under localized defects via a compact coupling argument (Theorem 4.7), a technique absent from the intrinsic-metric approach.
+Third, we establish a divergence criterion (Theorem 4.10) that guarantees essential self-adjointness even in the absence of $\chi$-completeness, thereby resolving in higher dimensions the open question raised in [BaGoJe] about the necessity of geometric completeness.
+Finally, our setting is tailored to clique complexes arising from weighted graphs, where cut-off functions are defined on vertices and extended by averaging a natural framework for applications in topological data analysis which allows for explicit energy estimates and concrete counterexamples (Section 5) that illustrate the sharpness of our results.
 In this paper, we have established a comprehensive framework for the essential self-adjointness of discrete Hodge Laplacians on higher-dimensional simplicial complexes, based on geometric completeness conditions adapted to the simplicial structure.
@@
 Hence, $DF_{k}\to DF$ in $\mathcal{H}$, which means $F_{k}\to F$ in the graph norm of $D$. Therefore, $F\in\mathrm{Dom}(D_{\min})$, so $D_{\min}=D_{\max}$, i.e., $D$ is essentially self-adjoint.
-Finally, since $L=D^{2}$ and $D$ is essentially self-adjoint, it follows that $L$ is also essentially self-adjoint on $\bigoplus_{i=0}^{n}C_{c}^{i}(V)$.
-This completes the proof. ∎
+Finally, since $L=D^{2}$ and $D$ is essentially self-adjoint, it follows that $L$ is also essentially self-adjoint on $\bigoplus_{i=0}^{n}C_{c}^{i}(V)$. This completes the proof. ∎
 This result is particularly relevant in topological data analysis, where point-cloud data may yield incomplete or noisy simplicial complexes.
 By adapting the weights to local simplex reliability, one can regularize the spectrum and ensure a well-defined Hodge decomposition.
+Acknowledgements: The authors are grateful to Prof.
+P.
+Bartmann and Prof.
+M.
+Keller for their valuable comments and constructive suggestions, which have contributed significantly to improving this paper.